OpenStudy (anonymous):

1. You are pushing a 150-kg wooden crate in a straight line a distance of 4.5 m across a wooden floor. The crate is speeding up at 0.38 m/s2. The static and kinetic coefficients of friction are 0.42 and 0.30, respectively. What is the work done by friction on the crate?

OpenStudy (anonymous):

I actually need to know what is the work done by YOU on the crate?

OpenStudy (anonymous):

also

OpenStudy (anonymous):

The crate is moving, so you only need to use the kinetic coefficient of friction: Force of friction = Kinetic coefficient*mass*gravitational acceleration Force of friction = 0.30*150kg*9.81m/s^2 = 441.45 Newtons Work = Force*Distance = 441.45N*4.5m = 1986.525 Joules Total Force = Force applied by you - Force of friction Total Force = mass*acceleration = 150kg*0.38m/s^2 = 57 Newtons Force applied by you = 441.45 Newtons + 57 Newtons = 490.95 Newtons Work done by you = Force applied by you*Distance = (490.95N)(4.5m) = 2209.275 Joules

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

No probs..